# Unraveling Indirect In-Context Learning Using Influence Functions

### v2

@@ Abstract
In this work, we introduce a novel paradigm for generalized In-Context Learning (ICL), termed Indirect In-Context Learning. In Indirect ICL, we explore demonstration selection strategies tailored for two distinct real-world scenarios: Mixture of Tasks and Noisy ICL. We systematically evaluate the effectiveness of Influence Functions (IFs) as a selection tool for these settings, highlighting the potential of IFs to better capture the informativeness of examples within the demonstration pool. For the Mixture of Tasks setting, demonstrations are drawn from 28 diverse tasks, including MMLU, BigBench, StrategyQA, and CommonsenseQA. We demonstrate that combining BertScore-Recall (BSR) with an IF surrogate model can further improve performance, leading to average absolute accuracy gains of 0.37% and 1.45% for 3-shot and 5-shot setups when compared to traditional ICL metrics. In the Noisy ICL setting, we examine scenarios where demonstrations might be mislabeled or have adversarial noise. Our experiments show that reweighting traditional ICL selectors (BSR and Cosine Similarity) with IF-based selectors boosts accuracy by an average of 2.90% for Cosine Similarity and 2.94% for BSR on noisy GLUE benchmarks. For the adversarial sub-setting, we show the utility of using IFs for task-agnostic demonstration selection for backdoor attack mitigation. Showing a 32.89% reduction in Attack Success Rate compared to task-aware methods. In sum, we propose a robust framework for demonstration selection that generalizes beyond traditional ICL, offering valuable insights into the role of IFs for Indirect ICL.
@@ 1 Introduction
In-Context Learning (ICL) has emerged as a powerful method for utilizing large language models (LLMs) to handle novel tasks at inference Mann et al. (2020); Min et al. (2022). Unlike traditional approaches that require task-specific fine-tuning, ICL allows a single model to adapt to different tasks without additional training, relying solely on the demonstrations provided in the context. This flexibility not only reduces the cost of task adaptation but also offers a transparent and easily customizable way of guiding the model’s behavior Liu et al. (2021a); Wei et al. (2022). By leveraging the context provided in prompts, ICL has been shown to improve both generalization across diverse tasks and reasoning abilities Anil et al. (2022); Drozdov et al. (2022). Despite its advantages, the success of ICL is closely tied to the choice of demonstrations used in the prompt. Even slight variations in these demonstrations can significantly influence the model’s performance, as shown in numerous studies Zhao et al. (2021); Liu et al. (2021a); Lu et al. (2022).
Figure 1: Example showcasing demonstration selection for Indirect ICL using Influence Functions (IFs). Consider web corpora with many tasks (different from the end-task) and noisy data– Indirect ICL can be formalized as: Mixture of Tasks (Section 3.1) and Noisy (Section 3.2) ICL, respectively. In MoT, for a given target task (e.g. Medical Genetics), we first filter from this (indirect) pool of candidate demonstrations using BertScore and Cosine Similarity, then re-rank with IFs to select suitable demonstrations (e.g. High-School Biology). For Noisy ICL, we leverage IFs to filter out the Noisy ICL before conducting ICL with the remaining clean demonstrations.
@@
Traditional ICL makes numerous assumptions that restrict its applicability to real-world problem domains. For instance, traditional ICL Min et al. (2021); Conneau (2019); Halder et al. (2020) assumes that demonstrations to be selected are directly and accurately annotated for the end-task. However, this is not always the case – for low-resource, sparse, or specialized domains, end-task information and labeled demonstrations might not be available. [^1] Consider the cases where we need to utilize ICL for diagnosing rare medical conditions, niche programming languages or indigenous spoken languages. Similarly, when LLMs are deployed as services, the user query or the end task itself could be unknown beforehand, let alone providing direct demonstrations at inference. [^2] Our proposed method can improve performance by selecting relevant demonstrations from a task agnostic pool of labeled data at test time. Thus, in this paper, we explore a more generalized setting for ICL, which we refer to as Indirect ICL.
@@
In Indirect ICL, we aim to provide indirect (or incidental) supervision Yin et al. (2023); Li et al. (2024) by selecting demonstrations from a pool of examples where the majority are not directly suited to the end task due to severe distribution/covariate shifts. This includes selecting demonstrations from a pool that predominantly consists of demonstrations belonging to other tasks, with few demonstrations from the end task possibly included. Additionally, the demonstration set may be mislabeled by humans Yan et al. (2014); Zhu et al. (2022) or LLMs Wu et al. (2023). Since the effectiveness of ICL heavily relies on the quality of demonstrations selected Kossen et al. (2024); Wu et al. (2022); Wang et al. (2024), selecting the most helpful indirect demonstrations becomes imperative in these situations.
@@
Despite these potential issues with the demonstration set, we wish to pave the way for extracting maximal benefit from any type of annotated dataset, irrespective of label purity or task relatedness. Thus, in order to combat the aforementioned issues with sub-optimal datasets for ICL, we leverage Influence Functions (IFs) Hampel (1974); Cook and Weisberg (1980). IFs offer a formal method for assessing how individual training data points affect model predictions. They have proven effective in a range of downstream machine learning tasks, including mislabeled data detection Koh and Liang (2017); Pruthi et al. (2020), optimal subset selection Feldman and Zhang (2020); Guo et al. (2020); Xia et al. (2024), model interpretation Han et al. (2020); Grosse et al. (2023); Chhabra et al. (2024b), data attribution Bae et al. (2024), data valuation Choe et al. (2024) and analyzing model biases Wang et al. (2019); Kong et al. (2021).
@@
Traditional (direct) ICL methods that use metrics such as BertScore-Recall (BSR; Gupta et al. 2023a) and cosine similarity Reimers (2019) inherently rely on the semantic similarity between demonstrations and test samples. In this paper, we posit that IFs can be a reasonable measure of affinity between the end task and any (indirect) demonstrations. We show that it is practical to use IFs to identify candidate demonstrations that represent a close inductive bias with the end-task, and utilize this information for highly accurate demonstration selection in the challenging Indirect ICL setting. As our experiments and results will demonstrate, this is indeed the case, and we find that IFs can aid in improved performance when simple semantic similarity is insufficient for demonstration selection. We provide additional examples of practical applications of Indirect ICL in Appendix A.
@@
In sum, our work advances ICL demonstration selection and makes the following key contributions and findings:
@@
  • For Mixture of Tasks, combining an IF Surrogate model with BertScore-Recall (BSR) can lead to a 0.37% and 1.45% average absolute increase in performance for $k=3$ and $k=5$ shots compared to the best performing traditional ICL metric.
  • For Noisy ICL, we observe that undertaking a weighted average selection using traditional ICL selectors (BSR and Cosine Similarity) and IF based selectors increases the absolute accuracy on average by 2.90% for Cosine and 2.94% for BSR for mislabeled samples. IFs can also lead to a 32.89% reduction in the Attack Success Rate for task-agnostic demonstration selection for backdoor attack mitigation.
@@ 2.1 Traditional In-Context Learning
Demonstration Selection. In this work we study the problem of selecting $k$ in-context examples from a pool of $N\gg k$ labeled candidates. This is often necessary due to context length limits and cost considerations (Rubin et al., 2021; Gupta et al., 2023a). Formally, the goal is to select a subset $S\subset\{(x_{i},y_{i})\}_{i=1}^{N}$ of size $k$ that maximizes the probability of generating the desired $y_{\text{test}}$ when the LLM is conditioned on $x_{\text{test}}$ and $S$. It is noteworthy that prior studies mainly consider a task-dependent ICL scenario and assume that candidate demonstrations all directly match the end task Min et al. (2021); Conneau (2019); Halder et al. (2020).
@@ 2.2 Indirect In-Context Learning
Noisy ICL. To further generalize the problem of Indirect ICL, we also consider noisy supervision that is likely existing in the pool of demonstrations. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ represent the training dataset, where $x_{i}\in X$ is the input and $y_{i}\in Y$ is the corresponding binary label. We randomly select a percentage of the data points from $D$ and flip their labels. Once the noisy dataset is generated, we use it for ICL. Formally, given the noisy set of input-output pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ and a test input $x_{\text{test}}$, the goal is to conditionally generate the test output $y_{\text{test}}$ based on the noisy training data.
@@ 2.3 Influence Functions
The IF evaluates the effect of individual training data points on the estimation of model parameters Hampel (1974); Cook and Weisberg (1980); Martin and Yohai (1986). It measures the rate at which parameter estimates change when a specific data point is up-weighted.
@@
The IF $I_{\theta^{*}}(x_{k},y_{k})$ on the empirical risk minimizer $\theta^{*}$ is generalized to assess its effect on prediction loss Koh and Liang (2017). Given a validation dataset $D^{\mathcal{V}}:=\{(x^{\mathcal{V}}_{i},y^{\mathcal{V}}_{i})\}_{i=1}^{m}$, the influence of $(x_{k},y_{k})$ on the validation loss is defined as:
$I(x_{k},y_{k})\coloneqq\Bigg{(}$ $\frac{1}{m}\sum_{i=1}^{m}\nabla_{\theta}\ell(y^{\mathcal{V}}_{i},f_{\theta}(x^{\mathcal{V}}_{i}))\bigg{|}_{\theta=\theta^{*}}\Bigg{)}^{\top}$
×$\times I_{\theta^{*}}(x_{k},y_{k}).$
@@
$I(x_{k},y_{k})=-\sum_{i=1}^{m}\Big{(}$ $\nabla_{\theta}\ell(y^{\mathcal{V}}_{i},f_{\theta}(x^{\mathcal{V}}_{i}))^{\top}$
$H(\theta^{*})^{-1}\nabla_{\theta}\ell(y_{k},f_{\theta}(x_{k}))\Big{)}.$
@@
Remark. As discussed above, IFs assume convexity of the loss function, which does not hold for LLMs and deep neural networks. Even though the IF formulations we employ in this paper Kwon et al. (2023); Koh and Liang (2017) make this underlying assumption, we find through empirical observations that for indirect ICL, they can work well. Circumventing the convexity assumption in IF is an ongoing area of research Grosse et al. (2023); Chhabra et al. (2024a) and our framework is flexible enough to accommodate any future IF variants.
@@ 3.1 Selecting within Mixture of Tasks
  • A surrogate-model based method, where a lightweight surrogate model such as RoBERTa Liu (2019) is fine-tuned on the candidate demonstrations to compute influence.
@@
Formally, for each validation example $(x_{\text{val}},y_{\text{val}})$, we compute the influence of each training example $(x_{i},y_{i})\in D_{\text{train}}$, where $D_{\text{train}}$ is the set of the training examples. The influence score $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ quantifies the effect of $(x_{i},y_{i})$ on the loss function evaluated at $(x_{\text{val}},y_{\text{val}})$ [^3] We use a negative log-likelihood function as the loss function. Using these computed influence values, we select the top $k$ samples with the highest IF scores.
We compare two versions of computing the IF after extracting the gradient, DataInf Kwon et al. (2023) and TracIn Pruthi et al. (2020). DataInf uses an easy-to-compute closed-form expression, leading to better computational and memory complexities than other IF methods, more details in Appendix B. TracIn traces how the loss on the test point changes during the training process simply using an inner product of training and validation set gradients. Since it does not compute the Hessian matrix, it is faster than DataInf, but at the cost of lower estimation performance.
Additionally, we compare the influence-only methods described above with strong ICL baselines BertScore-Recall (BSR; Gupta et al. 2023a) and Cosine Similarity Reimers (2019). [^4] We use the implementation from Gupta et al. (2023a, b). These methods excel at capturing semantic similarity between validation and training examples. We also compare with a performant sparse information retrieval baseline algorithm, BM25 Jones et al. (2000).
@@ 3.2 Selecting Noisy ICL
where $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ is the influence value and $S(x_{i},y_{i})$ represents either the BertScore Zhang et al. (2019) or Cosine Similarity for the training example $(x_{i},y_{i})$. The top $k$ examples with the highest combined scores are selected as demonstrations. We will refer to this approach as IF Averaging.
In this setting, we compute influence values using our surrogate model approach. In addition to using DataInf, we also conduct influence experiments using the LiSSA IF method which is a second-order method to compute the inverse Hessian vector product Agarwal et al. (2017); Koh and Liang (2017). Although LiSSA is generally computationally expensive Kwon et al. (2023), we prioritize it over TracIn owing to its greater performance in detecting mislabeled samples, as the computational overhead is incurred only once in this setting [^5] We use different algorithms since Noisy ICL computes IF scores using all validation samples, enabling detection of noisy training samples. In contrast, MoT requires sample-specific IF scores, necessitating dynamic, per-instance computation. .

### v3

@@ Abstract
In this work, we introduce a novel paradigm for generalized In-Context Learning (ICL), termed Indirect In-Context Learning. In Indirect ICL, we explore demonstration selection strategies tailored for two distinct real-world scenarios: Mixture of Tasks and Noisy ICL. We systematically evaluate the effectiveness of Influence Functions (IFs) as a selection tool for these settings, highlighting the potential of IFs to better capture the informativeness of examples within the demonstration pool. For the Mixture of Tasks setting, demonstrations are drawn from 28 diverse tasks, including MMLU, BigBench, StrategyQA, and CommonsenseQA. We demonstrate that combining BertScore-Recall (BSR) with an IF surrogate model can further improve performance, leading to average absolute accuracy gains in 3-shot and 5-shot setups when compared to traditional ICL metrics. In the Noisy ICL setting, we examine scenarios where demonstrations might be mislabeled or have adversarial noise. Our experiments show that reweighting traditional ICL selectors (BSR and Cosine Similarity) with IF-based selectors boosts accuracy on noisy GLUE benchmarks. For the adversarial sub-setting, we show the utility of using IFs for task-agnostic demonstration selection for backdoor attack mitigation. Showing a reduction in Attack Success Rate compared to task-aware methods. In sum, we propose a robust framework for demonstration selection that generalizes beyond traditional ICL, offering valuable insights into the role of IFs for Indirect ICL.
@@ 1 Introduction
In-Context Learning (ICL) has emerged as a powerful method for utilizing large language models (LLMs) to handle novel tasks at inference (Mann et al., 2020; Min et al., 2022). Unlike traditional approaches that require task-specific fine-tuning, ICL allows a single model to adapt to different tasks without additional training, relying solely on the demonstrations provided in the context. This flexibility not only reduces the cost of task adaptation but also offers a transparent and easily customizable way of guiding the model’s behavior (Liu et al., 2021a; Wei et al., 2022). By leveraging the context provided in prompts, ICL has been shown to improve both generalization across diverse tasks and reasoning abilities (Anil et al., 2022; Drozdov et al., 2022). Despite its advantages, the success of ICL is closely tied to the choice of demonstrations used in the prompt. Even slight variations in these demonstrations can significantly influence the model’s performance, as shown in numerous studies (Zhao et al., 2021; Liu et al., 2021a; Lu et al., 2022).
@@
Traditional ICL makes numerous assumptions that restrict its applicability to real-world problem domains. For instance, traditional ICL (Min et al., 2021; Conneau, 2019; Halder et al., 2020) assumes that demonstrations to be selected are directly and accurately annotated for the end-task. However, this is not always the case – for low-resource, sparse, or specialized domains, end-task information and labeled demonstrations might not be available. Similarly, when LLMs are deployed as services, the user query or the end task itself could be unknown beforehand, let alone providing direct demonstrations at inference. [^1] Our proposed method can improve performance by selecting relevant demonstrations from a task agnostic pool of labeled data at test time. Thus, in this paper, we explore a more generalized setting for ICL, which we refer to as Indirect ICL.
@@
Indirect ICL appears in real-world applications in several ways. For instance, it can be applied to diagnosing rare medical conditions, where no precedents or labeled demonstrations exist. Similarly, it is relevant for niche programming languages or indigenous spoken languages, which often lack sufficient labeled data. In Indirect ICL, we aim to provide indirect (or incidental) supervision (Yin et al., 2023; Li et al., 2024) by selecting demonstrations from a pool of examples where the majority are not directly suited to the end task due to severe distribution/covariate shifts. This includes selecting demonstrations from a pool that predominantly consists of demonstrations belonging to other tasks, with few demonstrations from the end task possibly included. Additionally, the demonstration set may be mislabeled by humans (Yan et al., 2014; Zhu et al., 2022) or LLMs (Wu et al., 2023). Since the effectiveness of ICL heavily relies on the quality of demonstrations selected (Kossen et al., 2024; Wu et al., 2022; Wang et al., 2024), selecting the most helpful indirect demonstrations becomes imperative in these situations. We provide detailed examples of practical applications of Indirect ICL in Appendix A.
Figure 1: Example showcasing demonstration selection for Indirect ICL using Influence Functions (IFs). Consider web corpora with many tasks (different from the end-task) and noisy data– Indirect ICL can be formalized as: Mixture of Tasks (Section˜3.1) and Noisy (Section˜3.2) ICL, respectively. In MoT, for a given target task (e.g. Medical Genetics), we first filter from this (indirect) pool of candidate demonstrations using BertScore and Cosine Similarity, then re-rank with IFs to select suitable demonstrations (e.g. High-School Biology). For Noisy ICL, we leverage IFs to filter out the Noisy ICL before conducting ICL with the remaining clean demonstrations.
Despite these potential issues with the demonstration set, we wish to pave the way for extracting maximal benefit from any type of annotated dataset, irrespective of label purity or task-relatedness. Moreover, existing approaches designed for Direct ICL often fail to generalize effectively to the Indirect ICL setting. For example, in an ablation study reported in Appendix C.7, we observe that standard retrieval methods such as Cosine Similarity and BERTScore-Recall select demonstrations from related tasks only 33.90% and 36.03% of the time, respectively, when evaluated on a pool where the vast majority of tasks are unrelated.
@@
Thus, in order to combat the aforementioned issues with sub-optimal datasets for ICL and sub-optimal demonstration selection strategies, we leverage Influence Functions (IFs) (Hampel, 1974; Cook & Weisberg, 1980). IFs offer a formal method for assessing how individual training data points affect model predictions. They have proven effective in a range of downstream data-centric learning tasks, including mislabeled data detection (Koh & Liang, 2017; Pruthi et al., 2020), optimal subset selection (Feldman & Zhang, 2020; Guo et al., 2020; Xia et al., 2024), model interpretation (Han et al., 2020; Grosse et al., 2023; Chhabra et al., 2024), data attribution (Bae et al., 2024), data valuation (Choe et al., 2024) and analyzing model biases (Wang et al., 2019; Kong et al., 2021).
@@
Traditional (direct) ICL methods that use metrics such as BertScore-Recall (BSR; Gupta et al. 2023a) and cosine similarity (Reimers, 2019) inherently rely on the semantic similarity between demonstrations and test samples. In this paper, we posit that IFs can be a reasonable measure of affinity between the end task and any (indirect) demonstrations. We show that it is practical to use IFs to identify candidate demonstrations that represent a close inductive bias with the end-task, and utilize this information for highly accurate demonstration selection in the challenging Indirect ICL setting. As our experiments and results will demonstrate, this is indeed the case, and we find that IFs can aid in improved performance when simple semantic similarity is insufficient for demonstration selection.
@@
Contributions. In sum, our work advances ICL demonstration selection and makes the following key contributions and findings:
@@
  • For Mixture of Tasks, combining an IF Surrogate model with BertScore-Recall (BSR) can lead to an increase in performance for $k=3$ and $k=5$ shots compared to the best performing traditional ICL metric.
  • For Noisy ICL, we observe that undertaking a weighted average selection using traditional ICL selectors (BSR and Cosine Similarity) and IF based selectors increases the absolute accuracy for mislabeled samples. IFs can also lead to a reduction in the Attack Success Rate for task-agnostic demonstration selection for backdoor attack mitigation.
@@ 2.1 Traditional In-Context Learning
Demonstration Selection. In this work we study the problem of selecting $k$ in-context examples from a pool of $N\gg k$ labeled candidates. This is often necessary due to context length limits and cost considerations (Rubin et al., 2021; Gupta et al., 2023a). Formally, the goal is to select a subset $S\subset\{(x_{i},y_{i})\}_{i=1}^{N}$ of size $k$ that maximizes the probability of generating the desired $y_{\text{test}}$ when the LLM is conditioned on $x_{\text{test}}$ and $S$. It is noteworthy that prior studies mainly consider a task-dependent ICL scenario and assume that candidate demonstrations all directly match the end task (Min et al., 2021; Conneau, 2019; Halder et al., 2020).
@@ 2.2 Indirect In-Context Learning
Noisy ICL. To further generalize the problem of Indirect ICL, we also consider noisy supervision that is likely existing in the pool of demonstrations. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote the training dataset, where $x_{i}\in X$ is the input and $y_{i}\in Y$ is the corresponding binary label. We randomly select a percentage of data points from $D$ and flip their labels. Once the noisy dataset is generated, we use it for ICL. Given the noisy set of input-output pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ and a test input $x_{\text{test}}$, the goal is to conditionally generate the test output $y_{\text{test}}$ based on the noisy training data.
@@ 2.3 Influence Functions
The IF evaluates the effect of individual training data points on the estimation of model parameters (Hampel, 1974; Cook & Weisberg, 1980; Martin & Yohai, 1986). It measures the rate at which parameter estimates change when a specific data point is up-weighted.
@@
The IF $I_{\theta^{*}}(x_{k},y_{k})$ on the empirical risk minimizer $\theta^{*}$ is generalized to assess its effect on prediction loss (Koh & Liang, 2017). Given a validation dataset $D^{\mathcal{V}}:=\{(x^{\mathcal{V}}_{i},y^{\mathcal{V}}_{i})\}_{i=1}^{m}$, the influence of $(x_{k},y_{k})$ on the validation loss is defined as:
$I(x_{k},y_{k})\coloneqq\Bigg($ $\frac{1}{m}\sum_{i=1}^{m}\nabla_{\theta}\ell(y^{\mathcal{V}}_{i},f_{\theta}(x^{\mathcal{V}}_{i}))\bigg|_{\theta=\theta^{*}}\Bigg)^{\top}\times I_{\theta^{*}}(x_{k},y_{k}).$
@@
$I(x_{k},y_{k})=-\sum_{i=1}^{m}\Big($ $\nabla_{\theta}\ell(y^{\mathcal{V}}_{i},f_{\theta}(x^{\mathcal{V}}_{i}))^{\top}H(\theta^{*})^{-1}\nabla_{\theta}\ell(y_{k},f_{\theta}(x_{k}))\Big).$
@@
Remark. As discussed above, IFs assume convexity of the loss function, which does not hold for LLMs and deep neural networks. Even though the IF formulations we employ in this paper (Kwon et al., 2023; Koh & Liang, 2017) make this underlying assumption, we find through empirical observations that for indirect ICL, they can work well. Circumventing the convexity assumption in IF is an ongoing area of research (Grosse et al., 2023; Chhabra et al., 2025) and our framework is flexible enough to accommodate any future IF variants.
@@ 3.1 Selecting within Mixture of Tasks
  • A surrogate-model based method, where a lightweight surrogate model such as RoBERTa (Liu, 2019) is fine-tuned on the candidate demonstrations to compute influence.
@@
Formally, for each validation example $(x_{\text{val}},y_{\text{val}})$, we compute the influence of each training example $(x_{i},y_{i})\in D_{\text{train}}$, where $D_{\text{train}}$ is the set of the training examples. The influence score $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ quantifies the effect of $(x_{i},y_{i})$ on the loss function evaluated at $(x_{\text{val}},y_{\text{val}})$ [^2] We use the standard negative log-likelihood as the loss function.. Using these computed influence values, we select the top $k$ samples with the highest IF scores.
We compare two versions of computing the IF after extracting the gradient, DataInf (Kwon et al., 2023) and TracIn (Pruthi et al., 2020). DataInf uses an easy-to-compute closed-form expression, leading to better computational and memory complexities than other IF methods, more details in Appendix B. TracIn traces how the loss on the test point changes during the training process simply using an inner product of training and validation set gradients. Since it does not compute the Hessian matrix, it is faster than DataInf, but at the cost of lower estimation performance.
Additionally, we compare the influence-only methods described above with strong ICL baselines BertScore-Recall (BSR; Gupta et al. 2023a) and Cosine Similarity (Reimers, 2019). These methods excel at capturing semantic similarity between validation and training examples. We also compare with a performant sparse information retrieval baseline algorithm, BM25 (Jones et al., 2000).
@@ 3.2 Selecting Noisy ICL
where $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ is the influence value and $S(x_{i},y_{i})$ represents either the BertScore (Zhang et al., 2019) or Cosine Similarity for the training example $(x_{i},y_{i})$. The top $k$ examples with the highest combined scores are selected as demonstrations. We will refer to this approach as IF Averaging.
In this setting, we compute influence values using our surrogate model approach. In addition to using DataInf, we also conduct influence experiments using the LiSSA IF method which is a second-order method to compute the inverse Hessian vector product (Agarwal et al., 2017; Koh & Liang, 2017). Although LiSSA is generally computationally expensive (Kwon et al., 2023), we prioritize it over TracIn owing to its greater performance in detecting mislabeled samples, as the computational overhead is incurred only once in this setting [^3] We use different influence algorithms since Noisy ICL computes IF scores using all validation samples, enabling detection of noisy training samples. In contrast, MoT requires sample-specific IF scores, necessitating dynamic and per-instance computation. .